A Bibliography of The ACM Journal of Computer Documentation

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

10 June 2016 Version 1.08

Title word cross-reference

1991-1994 [16].

Academic [6]. Academics [146]. Achieving [4]. ACM [108]. Acquisition [18]. Adoption [86]. age [187]. Aided [10]. Ain’t [95]. amateurs [198]. Amid [4]. Analysis [149, 77, 19, 105]. Analytical [27, 26, 47]. Annotated [52]. Anticipations [127]. application [202]. Applications [156]. Approach [72, 128, 164, 199]. Approaches [128, 134]. Aristotle [5]. Article [8]. Artistic [197, 198, 200, 196]. Assessing [211]. Assessment [93]. Assisting [135]. Assumptions [82]. asynchronous [174]. Atlas [93, 92, 94]. Atoms [129]. Audience [105, 167]. Author [36, 130]. Author-Reader [36]. Authority [110]. authors [202].

Backward [115]. Balance [111]. Based [149, 61, 64, 71]. Belong [114]. Bernadette [181, 180, 182]. Between [92, 202]. Beyond [143]. Bibliography [139, 52]. Biomedical [148]. Bits [129]. Blind [167]. Book [37]. Book-Review [37]. Books [42, 43]. Both [146]. Bounds [137]. Brasseur [7, 8]. Business [73, 168].

Call [140, 108, 53]. Can [107, 19]. Can’t [68]. capital [182]. Case [112, 113, 174, 214]. Cassini [112, 113]. Catalyst [45]. Categorical [134]. Centered [166]. century [218]. Challenge [146]. Change [9]. Changes [92]. changing [202]. Clark [2, 5]. Classic [124, 157, 195]. Cognition [80, 96]. Cognitive [46]. Cognitivist [143]. Coin [181, 180, 182]. collaboration [174]. College [25]. Commentaries [87, 37, 130, 124, 213, 152, 180, 157, 195, 200, 206]. Commentary [7, 34, 62, 217, 32, 191, 197, 112, 17, 144, 101, 145, 39, 113, 31, 38, 33, 102, 94, 203, 204, 171, 188, 205, 198]. Comments [8, 212]. common [213]. Communication [2, 1, 4, 67, 104, 155, 5, 114, 218, 25, 73]. communicator [217, 215]. Communicators [158, 121]. Community [75, 106]. Comparative [32]. Comparing [128]. Comparisons [84]. Complex [50, 51]. complexity [174]. Composition [25]. Compromised [3]. Computer [9, 36, 10, 159, 163, 64, 71, 68, 82, 192]. Computer-Aided [10]. Computer-Based [64, 71]. Computer-Mediated [82]. Computing [147, 151]. Concepts [80]. Concerns [60]. conditions [202]. Conference [90, 53, 58]. Confessions [154]. Conflict [4]. Consciousness [80]. Consequences [199]. Constructing [164, 96]. Content [61]. Content-Based [61]. Contents [25, 48, 46, 57, 73, 97]. Context [63, 167, 186, 194]. Contextual [30, 29, 31]. Continuing [87]. Conversation [69, 66, 132, 181, 171]. Conversations [194]. CoRR [149, 147, 152]. Correct [2]. Counterfeit [182]. Credibility [150]. Cross [85]. Cross-Cultural [85]. Cultural [77, 85]. Culture [83]. Current [82]. Cyberspace [105].

Declaration [35]. Deep [168]. Defining [105]. Departments [114]. Design [9, 10, 21, 89, 79, 23, 203, 204, 209, 205, 206, 202]. Designer [122]. Designers [77, 202, 162]. Designing [103, 99]. Details [65]. developers [202]. Development [75]. Different [11, 107]. Dilemma [150].

Dimensional [72]. Discipline

[197, 198, 200, 196]. discusses [160].

disputes [202]. distributed [214]. Diverse [78]. Do [68, 114]. Document [21, 122]. Documentation [28, 36, 49, 12, 32, 159, 163, 133, 164, 47, 14, 132, 134, 192, 97]. Documenting [51]. Documents [56, 211]. Doheny [70]. Driven [28]. Duffy [27, 26].

Easier [91]. Ecologies

[154, 155, 156, 164, 153]. Edit [91, 93, 94, 95, 92]. Editorial [172, 175, 201, 141]. Edmond [198]. Educating [51]. Education [118, 214]. Edward [87, 84]. Egoless [197, 198, 196, 200]. Emergence [166, 210]. Encoding [20]. Engineer [122]. Engineering [83, 197, 198, 200, 199, 196]. Engineers [107]. English [136, 137]. Entering [13]. Era [13, 14, 189]. eracies [137]. ergonomics [173]. Ethical [5]. Ethics [2, 1, 6, 3]. Ethnography [79]. evaluation [173]. even [160]. Everyday [67]. Evolution [21]. Ex [137]. Examples [20]. Expanding [136, 143, 214]. Expectations [170]. Expertise [110, 20]. Explanations [20, 84, 88]. explorations [218]. Exploring [167]. Extended [66]. extensions [216]. Extraordinary [104].

Facilitating [169]. Factors [45]. Farewell [15]. Farina [70]. Fifteen [120]. Fill [8]. Filling [10]. First [203, 38, 33, 151, 218]. Fit [168]. Flimsy [2]. Flower [96]. foraging [217, 215, 216]. Formulas [161, 160]. Forward [115]. Four [80]. Fragile [2]. Framework [143]. Frank [188]. fulfilled [141]. Further [63]. Future [12, 61, 45].

Gardener [154]. Generation [188, 184]. Genre [164]. Geography [83]. Going [55]. Grammar [56]. Graph [9]. Graphics [72, 99]. Graphing [9]. Gregory [2]. groups [174]. Groupware [85, 82]. GSS [83]. Guises [49].

H [198]. Hackos [47]. Hail [15]. Halasz [185, 186, 188]. Happen [107]. Haramundanis [34, 33]. HCI [139]. Heart [155, 153]. help [193, 27, 26]. History [181, 180, 126, 218]. Hitting [67]. Holistic [133]. Home [67, 114]. HTML [56]. Human [45]. Hypermedia [183, 195, 184]. Hyperphoric [56]. Hypertext [127, 188, 173, 189].

Icon [35]. If [95]. II [71]. Images [84]. Impact [93]. implications [216]. Implicit [104]. Improving [197, 198, 200, 196]. Impulse [197, 198, 200, 196]. Include [136]. incomplete [194]. Increase [19]. Increasing [24]. Independence [35]. Indexing [55]. Indices [16]. Industry [28, 146]. Influence [170, 87]. Information [166, 32, 154, 158, 155, 103, 156, 109, 99, 208, 213, 217, 215, 216, 153, 209, 212, 162, 213]. Information-Supported [99]. informed [192]. Inquiry [30, 29, 31]. Instruction [64, 71]. Instructional [57]. Integrating [146]. Intentional [142, 144, 145]. Intentionality [144]. Interaction [132]. Interface [134, 209]. Interfaces [191, 89, 190]. Internationalization [52]. Interview [70]. Introduction [180, 37, 74, 40, 124, 157, 207, 195]. invisibilities [193]. Invisibility [155]. Invisible [105]. IPCC [140]. IPCC/SIGDOC [140]. Issue [74]. Issues [148, 189, 144, 188, 82, 185, 186, 184].

JCD [16]. Journal [73, 97]. Just [11].

Klare [160, 162]. Know [77]. Knowledge [18, 107, 109].

landscape [214]. Language [48]. Learn [117, 58]. Learners [116]. Learning [118, 142, 144, 145]. Light [126]. limitations [160]. Linda [96]. Linguistics [56]. lining [182]. Links [56]. Lit [137]. Literacies [138]. Literary [127]. Literature [32]. Little [191, 190, 193]. Localization [52]. Lodging [65]. Longo [181, 180, 182]. Look [24]. Looking [120, 115, 94].

Machine [127]. Machines [191, 193, 190]. Mainstream [61]. Make [79]. Making [91]. Management [181, 180]. Managing [47]. Manuals [91]. Map [100]. Mapping [214]. Margaret [144]. Markup [61]. Markups [56]. Marshall [94]. Martinez [144]. Mayer [22]. Meaning [87]. Measurement [158]. Mediated [82]. Meets [61, 187]. Mehlenbacher [27, 26]. Mental [23]. Merits [60]. Metaphor [170, 89]. Metaphors [170]. metaphysics [212]. Millennium [161, 183, 195]. Millwrights [171, 218]. Minimalism [120, 121]. Model [60]. Models [23]. Modularity [128]. Multimedia [66].

Nardi [153]. narrative [173]. NASA [112, 113]. Need [77, 202]. Needs [10]. Neighborhood [69, 66]. Networks [68]. neurotic [198]. News [98, 54, 81, 131]. Next [50, 188, 184]. Nonsignificant [144]. Norms [5]. NoteCards [188, 185, 184, 187]. Notes [192]. novice [217, 215].

O’Day [153]. OHCO [60]. Online [69, 148, 27, 169, 106, 104, 26, 174]. Only [75]. Open [164]. Open-System [164]. Organization [166, 125]. Other [144, 97]. Our [193]. Outlining [126]. Oxymoron [106].

Pages [42, 43]. Pages/Pages [43]. Palmer [27, 26]. panding [137]. Papers [140, 108, 53]. Part [64, 71]. Participating [169]. Past [45]. Pattern [8]. paving [209]. pedagogy [192]. perceptions [202]. perfect [187]. Performance [19, 86, 64, 71, 99]. Perspective [148, 1, 6, 22, 5]. Perspectives [88, 85, 78, 82]. Pharmaceutical [28]. Places [44]. plan [141]. Planning [217, 215, 216]. Plato [5]. Politics [165]. Possibilities [104]. Post [11, 13, 14]. Post-Usability [11, 13, 14]. Potential [121]. Power [111, 109]. Practical [41, 40, 39, 38]. Practice [75, 170, 22, 187]. Present [45]. Presentations [17]. Preview [123]. Principled [4]. Principles [84]. print [209]. Problem [19]. Problem-Solving [19]. Problems [49, 50, 82, 151]. Process [9, 165, 22, 167, 96]. Processes [51]. Processing [61]. Producing [208, 213, 212]. Product [93, 165]. Production [107]. Products [167]. Profession [54, 210]. Professional [83, 6, 202]. Profit [165, 167]. Program [98, 131, 123]. Programs [9]. Project [86, 112, 113]. Projects [169, 47]. Promises [121]. Prose [19]. Proven [121]. Publications [125]. publishing [141]. Purpose [167]. Pushing [100].

Quality [213, 197, 209, 198, 200, 208, 211, 196, 212].

R [87]. Readability [159, 158, 160, 161]. Readable [163]. Reader [36, 119]. readers [202]. Reading [23]. Real [75]. Realities [168]. Reality [42, 106]. Really [63, 59, 62, 13]. Recurring [49]. redesign [173]. Reflecting [92]. Reflections [184, 189, 188, 134, 210]. regime [203, 204, 205, 206, 202]. Registration [98, 131, 65]. Relationship [36]. reminiscences [178, 179, 177, 176]. Replacing [197, 198, 200, 196]. Reply [200, 206]. Repositories [148]. Repository [147, 151]. Representation [18, 29]. Reprint [124, 157, 195]. Research [148, 147, 151]. Response [213, 30, 130, 152].

Rethinking [36, 134]. Retreat [14].

Review [27, 154, 37, 26, 156, 47]. Revisited [95]. Rhetoric [3, 129]. Rhetorical [1, 5]. Rich [95]. Role [133, 185]. Room [76, 79].

Satisfaction [86]. Science

[181, 32, 180, 19, 46, 57]. Sciences [148]. Scientific [24]. Scope [99]. searching [182]. Second [34, 204, 24, 39]. Seeking [12]. Selecting [10]. Self [75]. sense [213]. Sequential [125]. Service [25, 48, 46, 57, 73, 97]. Seven [72, 189, 188, 185, 186, 184]. Seven-Dimensional [72]. SGML} [41, 40, 40, 39, 38]. shipwrights [218, 171]. Should [117]. Sidelines [169]. Sides [146]. SIG [81]. SIGDOC [178, 179, 140, 177, 139, 176, 108, 123]. SIGDOC94 [17]. SIGDOC97 [53, 58, 65]. SIGDOC98 [90, 98]. SIGDOC99 [131]. silver [182]. Simplifying [51]. Since [21]. Smart [95]. Social [96, 216, 202]. socially [192]. Software [83, 12, 165]. Solve [151]. Solving [19]. Some [210, 194]. Spaces [44]. Speak [107]. Speaks [32]. Speed [150]. Spot [167]. Spurious [180, 182, 181]. States [218]. Step [50]. Still [94]. STOP [127, 128, 126, 125, 129]. Strategies [23]. Structural [19]. Structured [128]. Students [117]. Studies [136]. Study [112, 113, 174]. Supported [99]. Supporting [118, 116, 119]. Sweet [114]. Synthesizing [78]. System [164]. Systems [188, 64, 71, 184, 183, 195].

Table [25, 48, 46, 57, 73, 97]. Take [2]. Taking [137]. Tapestry [84]. Task [51]. Tasks [29]. Teams [85, 86]. Teamwork [82]. Technical [2, 1, 110, 213, 181, 73, 180, 121, 107, 122, 155, 109, 5, 129, 114, 218, 208, 217, 215, 209, 199, 212]. Techniques [18]. technological [202]. Technologies [67]. Technology [28, 166, 86, 64, 71, 153]. Telementoring [169]. testing [173]. Text [63, 59, 62, 24, 60, 23]. Thanks [54]. their [217, 193, 215]. Thematic [125]. Theme [58]. theories [217, 215, 216]. Theory [75, 170, 22]. Thick [22]. Thin [22]. Thinking [92]. Third [62, 205]. thoughts [194]. Threats [104]. Too [54]. tool [173]. Travel [98, 131]. Trouble [94]. Tufte [87, 84]. twenty [218]. twenty-first [218]. Two [128].

Understandability [24]. Understanding [191, 164, 190]. United [218]. Unveiling [104]. Update [90]. Usability [11, 54, 165, 13, 92, 14, 168, 173, 214]. usable [209]. Usage [149]. Use [91, 118, 161]. Useful [158, 162]. User [91, 10, 93, 89, 135, 119, 51, 14, 94, 95, 92]. Users [116, 191, 103, 190]. Using [30, 31, 153].

value [217, 215]. Venture [165]. Venue [58]. View [76, 96]. Virtual [42, 135, 181]. Visible [48]. Visual [88, 87, 84]. Voices [80]. vs [150].

way [209]. Ways [120]. Web [42, 185, 189, 209, 187, 162]. Weiss [198]. We’re [95]. Where [114]. Wired [69, 66]. Without [43, 44]. Worker [166]. Worker-Centered [166]. Workplace [136]. World [142, 144, 145]. Worlds [107]. Worries [138]. Wright [101, 102]. Writer [122, 129]. Writers [107]. Writing [136, 110, 197, 128, 180, 198, 109, 200, 96, 199, 196, 181]. WWW [54].

XML [203, 204, 205, 206, 202].

References

Clark:1994:ETC

[1] Gregory Clark. Ethics in technical communication: A rhetorical perspective.

Asterisk: the journal of computer documentation, 18(3):5–10, August 1994. CODEN ASTRF7. ISSN 0731-1001.

Brockmann:1994:GCT

[2] R. John Brockmann. Gregory Clark’s take on technical communication ethics: Flimsy, fragile, but correct. Asterisk: the journal of computer documentation, 18(3):11–15, August 1994. CODEN AS-TRF7. ISSN 0731-1001.

Davis:1994:REC

[3] Michael Davis. The rhetoric of ethics compromised. Asterisk: the journal of computer documentation, 18(3):15– 19, August 1994. CODEN ASTRF7. ISSN 0731-1001.

Girill:1994:APC

[4] T. R. Girill. Achieving principled communication amid conflict. Asterisk: the journal of computer documentation, 18 (3):19–24, August 1994. CODEN AS-TRF7. ISSN 0731-1001.

Walzer:1994:ENT

[5] Arthur E. Walzer. Ethical norms for technical communication: Plato, Aristotle, and Clark’s rhetorical perspective. Asterisk: the journal of computer documentation, 18(3):25–32, August 1994. CODEN ASTRF7. ISSN 0731-1001.

Clark:1994:PEA

[6] Gregory Clark. Professional ethics from an academic perspective. Asterisk: the journal of computer documentation, 18 (3):32–38, August 1994. CODEN AS-TRF7. ISSN 0731-1001.

Allmendinger:1994:CB

[7] Leif Allmendinger. Commentary on Brasseur. Asterisk: the journal of computer documentation, 18(4):??, November 1994. CODEN ASTRF7. ISSN 0731-1001.

Borland:1994:CBF

[8] Russell Borland. Comments on Brasseur’s fill pattern article. Asterisk: the journal of computer documentation, 18(4):??, November 1994. CODEN AS-TRF7. ISSN 0731-1001.

Brasseur:1994:HCG

[9] Lee Brasseur. How computer graphing programs change the graph design process. Asterisk: the journal of computer documentation, 18(4):??, November 1994. CODEN ASTRF7. ISSN 0731-1001.

Gustafson:1994:CAD

[10] Jolene Gustafson. Computer-aided design: Selecting and filling users’ needs. Asterisk: the journal of computer documentation, 18(4):??, November 1994. CODEN ASTRF7. ISSN 0731-1001.

Brockmann:1995:PUJ

[11] R. John Brockmann. Not post-usability, just different usability. Asterisk: the journal of computer documentation, 19 (1):??, March 1995. CODEN ASTRF7. ISSN 0731-1001.

Farkas:1995:SFS

[12] David K. Farkas. Seeking the future of software documentation. Asterisk: the journal of computer documentation, 19 (1):??, March 1995. CODEN ASTRF7. ISSN 0731-1001.

Redish:1995:WRE

[13] Janice Redish. Are we really entering a post-usability era. Asterisk: the journal of computer documentation, 19 (1):??, March 1995. CODEN ASTRF7. ISSN 0731-1001.

Weiss:1995:RUU

[14] Edmund H. Weiss. The retreat from usability: User documentation in the post-usability era. Asterisk: the journal of computer documentation, 19(1): ??, March 1995. CODEN ASTRF7. ISSN 0731-1001.

Borland:1995:FH

[15] Russell Borland. Farewell and hail. Asterisk: the journal of computer documentation, 19(2):1–4, June 1995. CO-DEN ASTRF7. ISSN 0731-1001.

Borland:1995:IJ

[16] Russell Borland. Indices to JCD 1991-1994. Asterisk: the journal of computer documentation, 19(2):5–8, June 1995. CODEN ASTRF7. ISSN 0731-1001.

Mehlenbacher:1995:CSP

[17] Brad Mehlenbacher. Commentary on SIGDOC94 presentations. Asterisk: the journal of computer documentation, 19 (2):15–21, June 1995. CODEN AS-TRF7. ISSN 0731-1001.

Gaines:1995:KAR

[18] Brian Gaines and Mildred Shaw. Knowledge acquisition and representation techniques. Asterisk: the journal of computer documentation, 19(2):23–36, June 1995. CODEN ASTRF7. ISSN 0731-1001.

Mayer:1995:SAS

[19] Richard E. Mayer. Structural analysis of science prose: Can we increase problem-solving performance? Asterisk: the journal of computer documentation, 19(3):3–26, August 1995. CODEN AS-TRF7. ISSN 0731-1001.

Girill:1995:EEE

[20] T. R. Girill. Expertise, examples, explanations, and encoding. Asterisk: the journal of computer documentation, 19 (3):27–30, August 1995. CODEN AS-TRF7. ISSN 0731-1001.

Killingsworth:1995:EDD

[21] M. Jimmie Killingsworth and Martin E. Rosenberg. The evolution of document design since 1985. Asterisk: the journal of computer documentation, 19(3): 31–35, August 1995. CODEN ASTRF7. ISSN 0731-1001.

Smith:1995:PPM

[22] Catherine Smith. A process perspective on mayer: Thin theory for thick practice. Asterisk: the journal of computer documentation, 19(3):36–38, August 1995. CODEN ASTRF7. ISSN 0731-1001.

Wright:1995:RSM

[23] Patricia Wright. Reading strategies, mental models, and text design. Asterisk: the journal of computer documentation, 19(3):38–45, August 1995. CODEN ASTRF7. ISSN 0731-1001.

Mayer:1995:SLI

[24] Richard E. Mayer. A second look at increasing the understandability of scientific text. Asterisk: the journal of computer documentation, 19(3):45–50, August 1995. CODEN ASTRF7. ISSN 0731-1001.

Girill:1995:TCS

[25] T. R. Girill. Table of contents service for College Composition and Communication. Asterisk: the journal of computer documentation, 19(4):3–8, November 1995. CODEN ASTRF7. ISSN 0731-1001.

Merrill:1995:ARD

[26] Cara K. Merrill. Analytical review of [duffy, palmer, mehlenbacher] Online Help. Asterisk: the journal of computer documentation, 19(4):3–8, November 1995. CODEN ASTRF7. ISSN 0731-1001.

Compton:1995:ARD

[27] Diane Compton. Analytical review of [duffy, palmer, mehlenbacher] Online Help. Asterisk: the journal of computer documentation, 19(4):8–12, November 1995. CODEN ASTRF7. ISSN 0731-1001.

Berhnardt:1995:TDD

[28] Stephen A. Berhnardt. Technology driven documentation in the pharmaceutical industry. Asterisk: the journal of computer documentation, 19(4): 13–18, November 1995. CODEN AS-TRF7. ISSN 0731-1001.

Mirel:1996:CIR

[29] Barbara Mirel. Contextual inquiry and the representation of tasks. Asterisk: the journal of computer documentation, 20 (1):14–21, February 1996. CODEN AS-TRF7. ISSN 0731-1001.

Kleimann:1996:RUC

[30] Susan Kleimann. Response to using contextual inquiry. Asterisk: the journal of computer documentation, 20(1):22– 24, February 1996. CODEN ASTRF7. ISSN 0731-1001.

Simpson:1996:CUC

[31] Mark Simpson. A commentary on using contextual inquiry. Asterisk: the journal of computer documentation, 20 (1):25–28, February 1996. CODEN AS-TRF7. ISSN 0731-1001.

Girill:1996:ISS

[32] T. R. Girill. Information science speaks to documentation: A comparative commentary on the recent literature. Asterisk: the journal of computer documentation, 20(1):29–31, February 1996. CO-DEN ASTRF7. ISSN 0731-1001.

Williams:1996:FCH

[33] Thomas R. Williams. First commentary on Haramundanis. Asterisk: the journal of computer documentation, 20(2):9– 15, May 1996. CODEN ASTRF7. ISSN 0731-1001.

Beardslee:1996:SCH

[34] Deborah Beardslee. Second commentary on Haramundanis. Asterisk: the journal of computer documentation, 20(2): 16–17, May 1996. CODEN ASTRF7. ISSN 0731-1001.

Pedell:1996:TDI

[35] Brian Pedell. Toward a declaration of icon independence. Asterisk: the journal of computer documentation, 20(2): 18–21, May 1996. CODEN ASTRF7. ISSN 0731-1001.

Coney:1996:RAR

[36] Mary B. Coney and Carl S. Chatfield. Rethinking the author-reader relationship in computer documentation. Asterisk: the journal of computer documentation, 20(2):22–28, May 1996. CODEN ASTRF7. ISSN 0731-1001.

Kades:1996:IBR

[37] Leigh Kades. Introduction to bookreview commentaries. Asterisk: the journal of computer documentation, 20 (2):29, May 1996. CODEN ASTRF7. ISSN 0731-1001.

Smith:1996:FCP

[38] Norman E. Smith. First commentary on Practical SGML. Asterisk: the journal of computer documentation, 20(2): 30–32, May 1, 1996. CODEN ASTRF7. ISSN 0731-1001.

Schick:1996:SCP

[39] William Schick. Second commentary on Practical SGML. Asterisk: the journal of computer documentation, 20(2): 33–35, May 1, 1996. CODEN ASTRF7. ISSN 0731-1001.

Price:1996:PSI

[40] Lynne A. Price. Practical SGML as an introduction to SGML. Asterisk: the journal of computer documentation, 20 (2):36–38, May 1, 1996. CODEN AS-TRF7. ISSN 0731-1001. Finds the Book’s strengths and weaknesses in its details.

Glushko:1996:HPP

[41] Robert J. Glushko. How practical is Practical SGML? Asterisk: the journal of computer documentation, 20(2): 39–43, May 1, 1996. CODEN ASTRF7. ISSN 0731-1001.

Dicks:1996:PBW

[42] R. Stanley Dicks. Pages, books, the Web, and virtual reality. Asterisk: the journal of computer documentation, 20 (3):9–13, August 1996. CODEN AS-TRF7. ISSN 0731-1001.

Kalmbach:1996:BPP

[43] James Kalmbach. Books without pages/pages without books. Asterisk: the journal of computer documentation, 20(3):14–17, August 1996. CODEN AS-TRF7. ISSN 0731-1001.

Mehlenbacher:1996:SP

[44] Brad Mehlenbacher. Spaces without places. Asterisk: the journal of computer documentation, 20(3):18–22, August 1996. CODEN ASTRF7. ISSN 0731-1001.

Ohnemus:1996:HFC

[45] Kenneth R. Ohnemus. Human(e) factors: A catalyst for the past, present, and future. Asterisk: the journal of computer documentation, 20(3):23–32, August 1996. CODEN ASTRF7. ISSN 0731-1001.

Girill:1996:TCSb

[46] T. R. Girill. Table of contents service for Cognitive Science. Asterisk: the journal of computer documentation, 20(3): 33–42, August 1996. CODEN ASTRF7. ISSN 0731-1001.

Storms:1996:ARH

[47] C. Gilbert Storms. Analytical review of [hackos] Managing Your Documentation Projects. Asterisk: the journal of computer documentation, 20(4):9– 13, November 1996. CODEN ASTRF7. ISSN 0731-1001.

Girill:1996:TCSa

[48] T. R. Girill. Table of contents service for Visible Language. Asterisk: the journal of computer documentation, 20(4): 14–25, November 1996. CODEN AS-TRF7. ISSN 0731-1001.

Draper:1997:NGR

[49] Stephen W. Draper. New guises for recurring problems in documentation. Asterisk: the journal of computer documentation, 21(1):15–18, February 1997. CODEN ASTRF7. ISSN 0731-1001.

Elser:1997:CPW

[50] Arthur G. Elser. Complex problems: What’s the next step? Asterisk: the journal of computer documentation, 21 (1):19–22, February 1997. CODEN AS-TRF7. ISSN 0731-1001.

Waite:1997:DCP

[51] Bob Waite. Documenting complex processes: Educating the user and simplifying the task. Asterisk: the journal of computer documentation, 21(1):23– 25, February 1997. CODEN ASTRF7. ISSN 0731-1001.

Uren:1997:ABI

[52] Emmanuel Uren. Annotated bibliography on internationalization and localization. Asterisk: the journal of computer documentation, 21(1):26–33, February 1997. CODEN ASTRF7. ISSN 0731-1001.

Smart:1997:SCC

[53] Karl Smart. SIGDOC97 conference call for papers. Asterisk: the journal of computer documentation, 21(1):34–35, February 1997. CODEN ASTRF7. ISSN 0731-1001.

Girill:1997:NPT

[54] T. R. Girill. News of the profession: Thanks and WWW usability too. Asterisk: the journal of computer documentation, 21(1):36–37, February 1997. CO-DEN ASTRF7. ISSN 0731-1001.

Mulvany:1997:WGI

[55] Nancy C. Mulvany. What’s going on in indexing? Asterisk: the journal of computer documentation, 21(2):10–15, May 1997. CODEN ASTRF7. ISSN 0731-1001.

Norton:1997:LLH

[56] David W. Norton. The linguistics of links: Hyperphoric grammar markups for HTML documents. Asterisk: the journal of computer documentation, 21 (2):16–21, May 1997. CODEN ASTRF7. ISSN 0731-1001.

Girill:1997:TCSa

[57] T. R. Girill. Table of contents service for Instructional Science. Asterisk: the journal of computer documentation, 21 (2):22–32, May 1997. CODEN ASTRF7. ISSN 0731-1001.

Smart:1997:SCT

[58] Karl Smart. SIGDOC97 conference: The theme, the venue, and how to learn more. Asterisk: the journal of computer documentation, 21(2):33–34, May 1997. CODEN ASTRF7. ISSN 0731-1001.

DeRose:1997:WTR

[59] Steven J. DeRose, David G. Durand, Elli Mylonas, and Allen H. Renear. What is text really? Asterisk: the journal of computer documentation, 21(3):1–25, August 1997. CODEN ASTRF7. ISSN 0731-1001.

Selber:1997:OMT

[60] Stuart A. Selber. The OHCO model of text: Merits and concerns. Asterisk: the journal of computer documentation, 21 (3):26–31, August 1997. CODEN AS-TRF7. ISSN 0731-1001.

Hill:1997:MMM

[61] Charles Hill. Markup meets the mainstream: The future of content-based processing. Asterisk: the journal of computer documentation, 21(3):32–35, August 1997. CODEN ASTRF7. ISSN 0731-1001.

Dicks:1997:TCQ

[62] R. Stanley Dicks. Third commentary on “What is Text Really?”. Asterisk: the journal of computer documentation, 21 (3):36–39, August 1997. CODEN AS-TRF7. ISSN 0731-1001.

DeRose:1997:FCQ

[63] Steven J. DeRose, David G. Durand, Elli Mylonas, and Allen H. Renear. Further context for “What is Text Really?”. Asterisk: the journal of computer documentation, 21(3):40–44, August 1997. CODEN ASTRF7. ISSN 0731-1001.

Reece:1997:PSTa

[64] Gloria A. Reece. Performance systems technology and computer-based instruction (Part I). Asterisk: the journal of computer documentation, 21(3):45– 55, August 1997. CODEN ASTRF7. ISSN 0731-1001.

Smart:1997:SLR

[65] Karl Smart. SIGDOC97: Lodging and registration details. Asterisk: the journal of computer documentation, 21(3): 56–63, August 1997. CODEN ASTRF7. ISSN 0731-1001.

Johnson:1997:WNE

[66] Bob Johnson. The Wired Neighborhood: an extended multimedia conversation. Asterisk: the journal of computer documentation, 21(4):1, November 1997. CODEN ASTRF7. ISSN 0731-1001.

Graham:1997:HHC

[67] Becky Graham. Hitting home: Communication technologies and the everyday.

Asterisk: the journal of computer documentation, 21(4):4–7, November 1997. CODEN ASTRF7. ISSN 0731-1001.

Rubingh:1997:WCN

[68] Mike Rubingh. What computer networks can’t do. Asterisk: the journal of computer documentation, 21(4):8– 10, November 1997. CODEN ASTRF7. ISSN 0731-1001.

Allen:1997:WNO

[69] Nancy Allen and Ann Blakeslee. The Wired Neighborhood: an online conversation. Asterisk: the journal of computer documentation, 21(4):11–15, November 1997. CODEN ASTRF7. ISSN 0731-1001.

Doheny-Farina:1997:ISD

[70] Stephen Doheny-Farina. Interview with stephen doheny-farina. Asterisk: the journal of computer documentation, 21 (4):16–19, November 1997. CODEN AS-TRF7. ISSN 0731-1001.

Reece:1997:PSTb

[71] Gloria A. Reece. Performance systems technology and computer-based instruction (Part II). Asterisk: the journal of computer documentation, 21(4):20– 25, November 1997. CODEN ASTRF7. ISSN 0731-1001.

Dowhal:1997:SDA

[72] Danny Dowhal. A seven-dimensional approach to graphics. Asterisk: the journal of computer documentation, 21(4): 26–37, November 1997. CODEN AS-TRF7. ISSN 0731-1001.

Girill:1997:TCSb

[73] T. R. Girill. Table of contents service for Journal of Business and Technical Communication. Asterisk: the journal of computer documentation, 21(4):38– 44, November 1997. CODEN ASTRF7. ISSN 0731-1001.

Mirel:1998:II

[74] Barbara Mirel. Introduction to this issue. Asterisk: the journal of computer documentation, 22(1):3–4, February 1998. CODEN ASTRF7. ISSN 0731-1001.

Bader:1998:WOS

[75] Gail Bader and James M. Nyce. When only the self is real: Theory and practice in the development community. Asterisk: the journal of computer documentation, 22(1):5–10, February 1998. CO-DEN ASTRF7. ISSN 0731-1001.

Bowker:1998:RV

[76] Geoffrey C. Bowker. A room with a view. Asterisk: the journal of computer documentation, 22(1):11–12, February 1998. CODEN ASTRF7. ISSN 0731-1001.

Dillon:1998:CAW

[77] Andrew Dillon. Cultural analysis and what designers need to know. Asterisk: the journal of computer documentation, 22(1):13–17, February 1998. CO-DEN ASTRF7. ISSN 0731-1001.

Rosson:1998:SDP

[78] Mary Beth Rosson. Synthesizing diverse perspectives. Asterisk: the journal of computer documentation, 22(1): 18–19, February 1998. CODEN AS-TRF7. ISSN 0731-1001.

Simonsen:1998:MRE

[79] Jesper Simonsen and Finn Kensing. Make room for ethnography in design!

Asterisk: the journal of computer documentation, 22(1):20–30, February 1998. CODEN ASTRF7. ISSN 0731-1001.

Nardi:1998:CCC

[80] Bonnie A. Nardi. Concepts of cognition and consciousness: Four voices. Asterisk: the journal of computer documentation, 22(1):31–48, February 1998. CO-DEN ASTRF7. ISSN 0731-1001.

Haramundanis:1998:NS

[81] Kathy Haramundanis. News of the SIG. Asterisk: the journal of computer documentation, 22(1):49–51, February 1998. CODEN ASTRF7. ISSN 0731-1001.

Vick:1998:PPC

[82] Rita Vick. Perspectives on and problems with computer-mediated teamwork: Current groupware issues and assumptions. Asterisk: the journal of computer documentation, 22(2):3–22, May 1998. CODEN ASTRF7. ISSN 0731-1001.

Auernheimer:1998:GPC

[83] Brent Auernheimer. GSS, professional culture, geography, and software engineering. Asterisk: the journal of computer documentation, 22(2):23–26, May 1998. CODEN ASTRF7. ISSN 0731-1001.

Needham:1998:ETV

[84] Russell K. Needham. Edward Tufte’s Visual Explanations: A tapestry of images, comparisons, and principles. Asterisk: the journal of computer documentation, 22(2):23–26, May 1998. CODEN AS-TRF7. ISSN 0731-1001.

Nordbotten:1998:PGC

[85] J. C. Nordbotten. Perspectives on groupware for cross-cultural teams. Asterisk: the journal of computer documentation, 22(2):27–29, May 1998. CO-DEN ASTRF7. ISSN 0731-1001.

R:1998:STA

[86] Panko, Raymond R. and Susan T. Kinney. Satisfaction, technology adoption, and performance in project teams. Asterisk: the journal of computer documentation, 22(2):30–33, May 1998. CO-DEN ASTRF7. ISSN 0731-1001.

Johnson:1998:VMC

[87] Robert. R. Johnson. Visual meaning: Commentaries on the continuing influence of edward R. tufte. Asterisk: the journal of computer documentation, 22 (2):34, May 1998. CODEN ASTRF7. ISSN 0731-1001.

Allen:1998:PVE

[88] Nancy Allen. Perspectives on Visual Explanations. Asterisk: the journal of computer documentation, 22(2):35–38, May 1998. CODEN ASTRF7. ISSN 0731-1001.

Marcus:1998:MDU

[89] Aaron Marcus. Metaphor design in user interfaces. Asterisk: the journal of computer documentation, 22(2):43–57, May 1998. CODEN ASTRF7. ISSN 0731-1001.

Bennett:1998:SCU

[90] Laurie Bennett. SIGDOC98 conference update. Asterisk: the journal of computer documentation, 22(2):58–60, May 1998. CODEN ASTRF7. ISSN 0731-1001.

Atlas:1998:UEM

[91] Marshall Atlas. The user edit: Making manuals easier to use. Asterisk: the journal of computer documentation, 22(3):5–6, August 1998. CODEN AS-TRF7. ISSN 0731-1001.

Schriver:1998:RAU

[92] Karen Schriver. Reflecting on Atlas’s User Edit: Changes in thinking about usability between 1981 and 1998. Asterisk: the journal of computer documentation, 22(3):7–11, August 1998. CODEN ASTRF7. ISSN 0731-1001.

Hayes:1998:AUE

[93] John R. Hayes. Atlas’s ‘The User Edit’: The impact on product assessment. Asterisk: the journal of computer documentation, 22(3):12–14, August 1998. CODEN ASTRF7. ISSN 0731-1001.

Wishbow:1998:SLT

[94] Nina Wishbow. Still looking for trouble: Commentary on Marshall Atlas’s ‘The User Edit’. Asterisk: the journal of computer documentation, 22(3):15–20, August 1998. CODEN ASTRF7. ISSN 0731-1001.

Atlas:1998:UER

[95] Marshall Atlas. The User Edit revisited, or if we’re so smart, why ain’t we rich? Asterisk: the journal of computer documentation, 22(3):24–26, August 1998. CODEN ASTRF7. ISSN 0731-1001.

Zimmerman:1998:LFS

[96] Beverly B. Zimmerman. Linda Flower and social cognition: Constructing a view of the writing process. Asterisk: the journal of computer documentation, 22(3):25–37, August 1998. CODEN AS-TRF7. ISSN 0731-1001.

Girill:1998:TCS

[97] T. R. Girill. Table of contents service for (the other) Journal of Documentation.

Asterisk: the journal of computer documentation, 22(3):38–63, August 1998. CODEN ASTRF7. ISSN 0731-1001.

Galt:1998:SPT

[98] Phyllis Galt. SIGDOC98 program, travel, and registration news. Asterisk: the journal of computer documentation, 22(3):64–70, August 1998. CODEN AS-TRF7. ISSN 0731-1001.

Wright:1998:DIS

[99] Patricia Wright. Designing informationsupported performance: The scope for graphics. Asterisk: the journal of computer documentation, 22(4):3–10, November 1998. CODEN ASTRF7. ISSN 0731-1001.

Borland:1998:PUM

[100] Russell Borland. Pushing us into the map. Asterisk: the journal of computer documentation, 22(4):11–15, November 1998. CODEN ASTRF7. ISSN 0731-1001.

Ramey:1998:CW

[101] Judith Ramey. Commentary on Wright. Asterisk: the journal of computer documentation, 22(4):16–20, November 1998. CODEN ASTRF7. ISSN 0731-1001.

Williams:1998:CW

[102] Thomas Williams. Commentary on Wright. Asterisk: the journal of computer documentation, 22(4):16–20, November 1998. CODEN ASTRF7. ISSN 0731-1001.

Rauch:1998:DIU

[103] Thyra Rauch. Designing information for users. Asterisk: the journal of computer documentation, 22(4):21–25, November 1998. CODEN ASTRF7. ISSN 0731-1001.

Jamieson:1998:UEP

[104] Lauren Jamieson. Unveiling the extraordinary possibilities and implicit threats of online communication. Asterisk: the journal of computer documentation, 22 (4):27–31, November 1998. CODEN AS-TRF7. ISSN 0731-1001.

Rosenberger:1998:AAC

[105] Lisa Rosenberger. Audience analysis in cyberspace: Defining the invisible. Asterisk: the journal of computer documentation, 22(4):32–36, November 1998. CODEN ASTRF7. ISSN 0731-1001.

Huse:1998:OCO

[106] Heidi Huse. Is ‘online community’ an oxymoron or a new reality? Asterisk: the journal of computer documentation, 22(4):37–44, November 1998. CODEN ASTRF7. ISSN 0731-1001.

Longo:1998:KPD

[107] Bernadette Longo. Knowledge production from different worlds: What can happen when technical writers speak for engineers. Asterisk: the journal of computer documentation, 22(4):45– 53, November 1998. CODEN ASTRF7. ISSN 0731-1001.

Selber:1998:CPA

[108] Stuart Selber. Call for papers for ACM SIGDOC 1999. Asterisk: the journal of computer documentation, 22(4):54– 56, November 1998. CODEN ASTRF7. ISSN 0731-1001.

Valentino:1999:ITW

[109] Mattio Valentino. Information, technical writing, knowledge, and power. Asterisk: the journal of computer documentation, 23(1):3–18, February 1999. CO-DEN ASTRF7. ISSN 0731-1001.

Clark:1999:TWA

[110] Gregory Clark. Technical writing and the authority of expertise. Asterisk: the journal of computer documentation, 23 (1):19–21, February 1999. CODEN AS-TRF7. ISSN 0731-1001.

Lundgren:1999:BP

[111] Regina Lundgren. Balance of power. Asterisk: the journal of computer documentation, 23(1):22–24, February 1999. CODEN ASTRF7. ISSN 0731-1001.

Kayten:1999:CCS

[112] Gerald Kayten. Commentary on a case study of NASA’s Cassini Project. Asterisk: the journal of computer documentation, 23(1):25–27, February 1999. CO-DEN ASTRF7. ISSN 0731-1001.

Shuman:1999:CCS

[113] Larry Shuman. Commentary on a case study of NASA’s Cassini Project. Asterisk: the journal of computer documentation, 23(1):25–27, February 1999. CO-DEN ASTRF7. ISSN 0731-1001.

Wishbow:1999:HSH

[114] Nina Wishbow. Home sweet home? where do technical communication departments belong. Asterisk: the journal of computer documentation, 23(1): 28–34, February 1999. CODEN AS-TRF7. ISSN 0731-1001.

Haramundanis:1999:LBL

[115] Kathy Haramundanis. Looking backward, looking forward. Asterisk: the journal of computer documentation, 23 (1):35–36, February 1999. CODEN AS-TRF7. ISSN 0731-1001.

Guzdial:1999:SLU

[116] Mark Guzdial. Supporting learners as users. Asterisk: the journal of computer documentation, 23(2):3–13, May 1999. CODEN ASTRF7. ISSN 0731-1001.

DiSessa:1999:HSS

[117] Andrea DiSessa. How should students learn? Asterisk: the journal of computer documentation, 23(2):14–18, May 1999. CODEN ASTRF7. ISSN 0731-1001.

Draper:1999:SUL

[118] Stephen Draper. Supporting use, learning, and education. (1999). Asterisk: the journal of computer documentation, 23 (2):19–24, May 1999. CODEN ASTRF7. ISSN 0731-1001.

VanderMeij:1999:SRU

[119] Hans Van der Meij. Supporting the reader as user. Asterisk: the journal of computer documentation, 23(2):25– 31, May 1999. CODEN ASTRF7. ISSN 0731-1001.

Dubinsky:1999:FWL

[120] James Dubinsky. Fifteen ways of looking at minimalism. Asterisk: the journal of computer documentation, 23(2): 34–47, May 1999. CODEN ASTRF7. ISSN 0731-1001.

Lodor:1999:PPP

[121] Eric Lodor. The proven and potential promises of minimalism for technical communicators. Asterisk: the journal of computer documentation, 23(2): 48–56, May 1999. CODEN ASTRF7. ISSN 0731-1001.

Marsh:1999:ETW

[122] C. Hugh Marsh. The engineer as technical writer and document designer. Asterisk: the journal of computer documentation, 23(2):57–61, May 1999. CO-DEN ASTRF7. ISSN 0731-1001.

Selber:1999:SPP

[123] Stuart Selber. SIGDOC 1999 program preview. Asterisk: the journal of computer documentation, 23(2):62–63, May 1999. CODEN ASTRF7. ISSN 0731-1001.

Waite:1999:ICR

[124] Bob Waite. Introduction to this classic reprint and commentaries. Asterisk: the journal of computer documentation, 23(3):2–3, August 1999. CODEN AS-TRF7. ISSN 0731-1001.

Tracey:1999:STO

[125] J. R. Tracey, D. E. Rugh, and W. S. Starkey. Sequential thematic organization of publications (STOP). Asterisk: the journal of computer documentation, 23(3):4–68, August 1999. CODEN AS-TRF7. ISSN 0731-1001.

Price:1999:SLH

[126] Jonathan Price. STOP: Light on the history of outlining. Asterisk: the journal of computer documentation, 23(3): 69–78, August 1999. CODEN ASTRF7. ISSN 0731-1001.

Bernstein:1999:AHS

[127] Mark Bernstein. Anticipations of hypertext: STOP and the literary machine. Asterisk: the journal of computer documentation, 23(3):79–86, August 1999. CODEN ASTRF7. ISSN 0731-1001.

Horn:1999:TAM

[128] Robert Horn. Two approaches to modularity: Comparing the STOP approach with structured writing. Asterisk: the journal of computer documentation, 23 (3):87–95, August 1999. CODEN AS-TRF7. ISSN 0731-1001.

Weiss:1999:BAT

[129] Edmund H. Weiss. Bits, atoms, and the technical writer: The rhetoric of STOP. Asterisk: the journal of computer documentation, 23(3):96–101, August 1999. CODEN ASTRF7. ISSN 0731-1001.

Starkey:1999:ARC

[130] Wendel S. Starkey. Author’s response to commentaries. Asterisk: the journal of computer documentation, 23(3):102– 103, August 1999. CODEN ASTRF7. ISSN 0731-1001.

Johnson-Eilola:1999:SPT

[131] Johndan Johnson-Eilola. SIGDOC99 program, travel, and registration news. Asterisk: the journal of computer documentation, 23(3):104–108, August 1999. CODEN ASTRF7. ISSN 0731-1001.

Winograd:1999:DIC

[132] Terry Winograd. Documentation, interaction, and conversation. Asterisk: the journal of computer documentation, 23 (4):3–7, November 1999. CODEN AS-TRF7. ISSN 0731-1001.

Quesenbery:1999:DHR

[133] Whitney Quesenbery. Documentation’s holistic role. Asterisk: the journal of computer documentation, 23(4):8–11, November 1999. CODEN ASTRF7. ISSN 0731-1001.

Wixon:1999:RDI

[134] Dennis Wixon. Rethinking documentation and interface: Reflections on categorical approaches. Asterisk: the journal of computer documentation, 23(4):

12–14, November 1999. CODEN AS-TRF7. ISSN 0731-1001.

Ober:1999:AVU

[135] John Ober. Assisting the virtual user. Asterisk: the journal of computer documentation, 23(4):15–21, November 1999. CODEN ASTRF7. ISSN 0731-1001.

Brady:1999:EES

[136] M. Ann Brady. Expanding English studies to include workplace writing. Asterisk: the journal of computer documentation, 23(4):23–26, November 1999. CO-DEN ASTRF7. ISSN 0731-1001.

Johnson:1999:EPL

[137] Evelyn Johnson. (ex)panding (lit)eracies: Taking English out of bounds. Asterisk: the journal of computer documentation, 23(4):27–29, November 1999. CODEN ASTRF7. ISSN 0731-1001.

Sullivan:1999:WAN

[138] Dale Sullivan. Worries about the new literacies. Asterisk: the journal of computer documentation, 23(4):30–34, November 1999. CODEN ASTRF7. ISSN 0731-1001.

Perlman:1999:HBS

[139] Gary Perlman. The HCI bibliography and SIGDOC. Asterisk: the journal of computer documentation, 23(4): 36, November 1999. CODEN ASTRF7. ISSN 0731-1001.

Jones:1999:ISC

[140] Susan B. Jones. IPCC/SIGDOC 2000 call for papers. Asterisk: the journal of computer documentation, 23(4):37– 38, November 1999. CODEN ASTRF7. ISSN 0731-1001.

Girill:2000:EPP

[141] T. R. Girill. Editorial: a publishing plan fulfilled. ACM Journal of Computer Documentation, 24(1):1–2, February 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-1/p1-girill/.

Martinez:2000:ILI

[142] Margaret Martinez. Intentional learning in an intentional world. ACM Journal of Computer Documentation, 24(1):3– 20, February 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-1/p3-martinez/.

Kirkley:2000:EBC

[143] Jamie Kirkley and Thomas Duffy. Expanding beyond a cognitivist framework. ACM Journal of Computer Documentation, 24(1):21–24, February 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-1/p21-kirkley/.

Mehlenbacher:2000:ION

[144] Brad Mehlenbacher. Intentionality and other “nonsignificant” issues in learning: Commentary on Margaret Martinez’s Intentional Learning in an Intentional World. ACM Journal of Computer Documentation, 24(1):25–30, February 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-1/p25-mehlenbacher/.

Russell:2000:CIL

[145] Thomas L. Russell. Commentary on Intentional Learning in an Intentional World. ACM Journal of Computer Documentation, 24(1):31–32, February 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/jcd/2000-
                                        24-1/p31-russell/.

Sutliff:2000:IAI

[146] Kristene Sutliff. Integrating academics and industry: A challenge for both sides. ACM Journal of Computer Documentation, 24(1):33–38, February 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/jcd/2000-
                                        24-3/p33-sutliff/.

Halpern:2000:CCR

[147] Joseph Y. Halpern. CoRR: A computing research repository. ACM Journal of Computer Documentation, 24(2): 41–48, May 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/jcd/2000-
                                        24-2/p41-halpern/.

Armbruster:2000:IOR

[148] David L. Armbruster. Issues of online research repositories from the perspective of the biomedical sciences. ACM Journal of Computer Documentation, 24(2): 49–53, May 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/jcd/2000-
                                        24-2/p49-armbruster/.

Carr:2000:UBA

[149] Les Carr, Steve Hitchcock, Wendy Hall, and Stevan Harnad. A usage based analysis of CoRR. ACM Journal of Computer Documentation, 24(2): 54–59, May 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-2/p54-carr/.

Prekeges:2000:DCV

[150] James Prekeges. The dilemma of credibility vs. speed. ACM Journal of Computer Documentation, 24(2):60– 63, May 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-2/p60-prekeges/.

vanLoon:2000:CRR

[151] A. J. van Loon. A computing research repository: Why not solve the problems first? ACM Journal of Computer Documentation, 24(2):64– 71, May 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-2/p64-van_loon/.

Halpern:2000:RCC

[152] Joseph Y. Halpern. A response to the commentaries on CoRR. ACM Journal of Computer Documentation, 24(2): 72–77, May 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-2/p72-halpern/.

Johnson:2000:NOI

[153] Robert R. Johnson. Nardi and O’Day’s information ecologies: using technology with heart. ACM Journal of Computer Documentation, 24(2):78, May 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-2/p78-johnson/.

Hart-Davidson:2000:CGR

[154] William Hart-Davidson. Confessions of a gardener: A review of information ecologies. ACM Journal of Computer Documentation, 24(2):79– 84, May 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-2/p79-hart-davidson/.

Ranney:2000:HIE

[155] Frances J. Ranney. At the heart of information ecologies: Invisibility and technical communication. ACM Journal of Computer Documentation, 24(2): 85–90, May 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-2/p85-ranney/.

Selfe:2000:RAI

[156] Dickie Selfe and Dawn Hayden. A review and applications of information ecologies. ACM Journal of Computer Documentation, 24(2):91–102, May 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-2/p91-selfe/.

Waite:2000:ICR

[157] Bob Waite. Introduction to this classic reprint and commentaries. ACM Journal of Computer Documentation, 24(3):105–106, August 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-3/p105-waite/.

Klare:2000:MRU

[158] George R. Klare. The measurement of readability: Useful information for communicators. ACM Journal of Computer Documentation, 24(3):107–121, August 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-3/p107-klare/.

Hargis:2000:RCD

[159] Gretchen Hargis. Readability and computer documentation. ACM Journal of Computer Documentation, 24(3):122– 131, August 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-3/p122-hargis/.

Redish:2000:RFE

[160] Janice Redish. Readability formulas have even more limitations than Klare discusses. ACM Journal of Computer Documentation, 24(3):132– 137, August 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-3/p132-redish/.

Schriver:2000:RFN

[161] Karen A. Schriver. Readability formulas in the new millennium: What’s the use? ACM Journal of Computer Documentation, 24(3):138–140, August 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-3/p138-schriver/.

Zibell:2000:KUI

[162] Kristin Zibell. Klare’s “useful information” is useful for Web designers. ACM Journal of Computer Documentation, 24(3):141–147, August 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-3/p141-zibell/.

Klare:2000:RCD

[163] George R. Klare. Readable computer documentation. ACM Journal of Computer Documentation, 24(3):148– 168, August 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-3/p148-klare/.

Spinuzzi:2000:GEO

[164] Clay Spinuzzi and Mark Zachry. Genre ecologies: An open-system approach to understanding and constructing documentation. ACM Journal of Computer Documentation, 24(3):169–181, August 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-3/p169-spinuzzi/.

Mirel:2000:PPP

[165] Barbara Mirel. Product, process, and profit: The politics of usability in a software venture. ACM Journal of Computer Documentation, 24(4):185–203, November 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-4/p185-mirel/.

Carlson:2000:ITE

[166] Patricia A. Carlson. Information technology and the emergence of a workercentered organization. ACM Journal of Computer Documentation, 24 (4):204–212, November 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). URL http://

www.acm.org/pubs/citations/journals/
jcd/2000-24-4/p204-carlson/.

Spinuzzi:2000:EBS

[167] Clay Spinuzzi. Exploring the blind spot: Audience, purpose, and context in “products, process, and profit”. ACM Journal of Computer Documentation, 24(4):213–219, November 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/jcd/2000-
24-4/p213-spinuzzi/.

Wiebe:2000:DRF

[168] Eric N. Wiebe. Deep realities:

The fit of usability in business. ACM Journal of Computer Documentation, 24(4):220–226, November 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

tronic). URL http://www.acm.org/

                                        pubs/citations/journals/jcd/2000-
                                        24-4/p220-wiebe/.

Harris:2000:PSO

[169] Judith B. Harris and Candace Figg. Participating from the sidelines, online: Facilitating telementoring projects. ACM Journal of Computer Documentation, 24(4):227–236, November 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/jcd/2000-
                                        24-4/p227-harris/.

Hamilton:2000:MTP

[170] Anne Hamilton. Metaphor in theory and practice: The influence of metaphors on expectations. ACM Journal of Computer Documentation, 24(4):237–253, November 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/jcd/2000-
                                        24-4/p237-hamilton/.

Jobst:2000:CCM

[171] Jack Jobst and Robert R. Johnson. A conversation and commentary on From Millwrights to Shipwrights. ACM Journal of Computer Documentation, 24(4):254–259, November 2000. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/jcd/2000-
                                        24-4/p254-jobst/.

Dicks:2001:Ea

[172] Stan Dicks. Editorial. ACM Journal of Computer Documentation, 25(1):1– 2, February 2001. CODEN AJCDBH.

ISSN 1527-6805 (print), 1557-9441 (electronic).

Gee:2001:EHN

[173] Kim Gee. The ergonomics of hypertext narrative: usability testing as a tool for evaluation and redesign. ACM Journal of Computer Documentation, 25(1):3– 16, February 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Chandler:2001:COG

[174] Hope E. Chandler. The complexity of online groups: a case study of asynchronous collaboration. ACM Journal of Computer Documentation, 25(1):17– 24, February 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Dicks:2001:Eb

[175] Stan Dicks. Editorial. ACM Journal of Computer Documentation, 25(2): 29, May 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Rigo:2001:SR

[176] Joe Rigo. SIGDOC reminiscences.

ACM Journal of Computer Documentation, 25(2):31–33, May 2001. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Patterson:2001:SR

[177] Diana Patterson. SIGDOC reminiscences 1981–88. ACM Journal of Computer Documentation, 25(2):34–39, May 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Brockman:2001:SR

[178] John Brockman. SIGDOC reminiscences. ACM Journal of Computer Documentation, 25(2):40–41, May 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Haramundanis:2001:SR

[179] Kathy Haramundanis. SIGDOC reminiscences. ACM Journal of Computer Documentation, 25(2):42–46, May 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Johnson:2001:ICS

[180] Bob Johnson. Introduction to commentaries on “Spurious Coin: A History of Science, Management, and Technical Writing” by Bernadette Longo. ACM Journal of Computer Documentation, 25 (2):47, May 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Gaskill:2001:VCB

[181] David Gaskill, Mary Been, Margaret Hundleby, and Pete Praetorius. A virtual conversation on Bernadette Longo’s Spurious Coin: A History of Science, Management, and Technical Writing. ACM Journal of Computer Documentation, 25(2):48–62, May 2001. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Trim:2001:CCS

[182] Michelle Trim. Counterfeit capital: searching for a silver lining in Bernadette Longo’s Spurious Coin. ACM Journal of Computer Documentation, 25(2):63–65, May 2001. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Waite:2001:HSNa

[183] Bob Waite. Hypermedia systems in the new millennium. ACM Journal of Computer Documentation, 25(3):69– 70, August 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Halasz:2001:RNS

[184] Frank G. Halasz. Reflections on Note-Cards: seven issues for the next generation of hypermedia systems. ACM Journal of Computer Documentation, 25(3): 71–87, August 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See comments [185, 186, 187, 188, 189].

Dobroth:2001:NWR

[185] Kate Dobroth. From NoteCards to the Web: the role of Halasz’s seven issues. ACM Journal of Computer Documentation, 25(3):88–91, August 2001. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [184].

Furuta:2001:HSI

[186] Richard Furuta. Halasz’s “seven issues” in context. ACM Journal of Computer Documentation, 25(3):92–95, August 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [184].

Marshall:2001:NAW

[187] Catherine C. Marshall. NoteCards in the age of the Web: practice meets perfect. ACM Journal of Computer Documentation, 25(3):96–103, August 2001. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [184].

Mylonas:2001:CFH

[188] Elli Mylonas. A commentary on Frank Halasz’s “Reflections on NoteCards: Seven Issues for the Next Generation of Hypertext Systems”. ACM Journal of Computer Documentation, 25(3):104– 108, August 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [184].

Halasz:2001:RSI

[189] Frank G. Halasz. Reflections on “Seven Issues”: hypertext in the era of the Web. ACM Journal of Computer Documentation, 25(3):109–114, August 2001. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [184].

Johnson-Eilola:2001:LMU

[190] Johndan Johnson-Eilola. Little machines: understanding users understanding interfaces. ACM Journal of Computer Documentation, 25(4):119–127, November 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See commentary [191].

Haramundanis:2001:CSM

[191] Kathy Haramundanis. Commentary on: “Little Machines: Understanding Users Understanding Interfaces”. ACM Journal of Computer Documentation, 25(4):128–131, November 2001. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [190].

Selber:2001:NTS

[192] Stuart A. Selber. Notes toward a socially informed pedagogy for computer documentation. ACM Journal of Computer Documentation, 25(4): 132–140, November 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Farkas:2001:OLH

[193] David K. Farkas. Our little help machines and their invisibilities. ACM Journal of Computer Documentation, 25(4):141–144, November 2001. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Johnson-Eilola:2001:CCS

[194] Johndan Johnson-Eilola. Conversations on context: some incomplete thoughts. ACM Journal of Computer Documentation, 25(4):145–147, November 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Waite:2001:HSNb

[195] Bob Waite. Hypermedia systems in the new millennium: introduction to this classic reprint and commentaries. ACM Journal of Computer Documentation, 25(4):148–152, November 2001. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Weiss:2002:EWI

[196] Edmond H. Weiss. Egoless writing:

improving quality by replacing artistic impulse with engineering discipline. ACM Journal of Computer Documentation, 26(1):3–10, February 2002. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See commentary [197, 198, 200].

Haramundanis:2002:CSW

[197] Kathy Haramundanis. Commentary on:

“Egoless Writing: Improving Quality by Replacing Artistic Impulse with Engineering Discipline”. ACM Journal of Computer Documentation, 26(1):11– 16, February 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [196].

Sauer:2002:WNA

[198] Geoffrey Sauer. We neurotic amateurs: a commentary on Edmond H. Weiss’s “Egoless Writing: Improving Quality by Replacing Artistic Impulse with Engineering Discipline”. ACM Journal of Computer Documentation, 26(1):17– 21, February 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [196].

Waite:2002:CEA

[199] Bob Waite. Consequences of the engineering approach to technical writing. ACM Journal of Computer Documentation, 26(1):22–26, February 2002. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Weiss:2002:RCS

[200] Edmond H. Weiss. Reply to commentaries on: “Egoless Writing: Improving Quality by Replacing Artistic Impulse with Engineering Discipline”. ACM Journal of Computer Documentation, 26(1):27–28, February 2002. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [196].

Dicks:2002:E

[201] Stan Dicks. Editorial. ACM Journal of Computer Documentation, 26(2): 31–32, May 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Wilkes:2002:XND

[202] Gilbert Vanburen Wilkes IV. XML and the new design regime: disputes between designers, application developers, authors, and readers in changing technological conditions and perceptions of social and professional need. ACM Journal of Computer Documentation, 26(2): 33–42, May 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See commentary [203, 204, 205, 206].

Burch:2002:FCS

[203] Glenda Burch. First commentary on “XML and the new design regime”. ACM Journal of Computer Documentation, 26(2):43–44, May 2002. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Haynes:2002:SCS

[204] Mir Haynes. Second commentary on “XML and the new design regime”. ACM Journal of Computer Documentation, 26(2):45–47, May 2002. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [202].

Payne:2002:TCS

[205] John Payne. Third commentary on “XML and the new design regime”. ACM Journal of Computer Documentation, 26(2):48–49, May 2002. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [202].

Wilkes:2002:RCS

[206] Gilbert Vanburen Wilkes IV. Reply to commentaries on “XML and the new design regime”. ACM Journal of Computer Documentation, 26(2):50–53, May 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Waite:2002:I

[207] Bob Waite. Introduction. ACM Journal of Computer Documentation, 26(3): 64–65, August 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Dean:2002:PQT

[208] Morris Dean. Producing quality technical information. ACM Journal of Computer Documentation, 26(3):66– 117, August 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See commentary [209, 212, 213].

Mandel:2002:QTI

[209] Theo Mandel. Quality technical information: paving the way for usable print and Web interface design. ACM Journal of Computer Documentation, 26(3):118– 125, August 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [208].

Grice:2002:SRE

[210] Roger A. Grice. Some reflections on the emergence of a profession. ACM Journal of Computer Documentation, 26(3):126– 129, August 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Smart:2002:AQD

[211] Karl L. Smart. Assessing quality documents. ACM Journal of Computer Documentation, 26(3):130–140, August 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Weiss:2002:MIQ

[212] Edmond H. Weiss. The metaphysics of information quality: comments on producing quality technical information. ACM Journal of Computer Documentation, 26(3):141–147, August 2002. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [208].

Dean:2002:RCP

[213] Morris Dean. Response to the commentaries on Producing Quality Technical Information: the common sense of producing quality technical information. ACM Journal of Computer Documentation, 26(3):148–154, August 2002. CO-DEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Grice:2002:MEL

[214] Roger Grice and Bill Hart-Davidson.

Mapping the expanding landscape of usability: the case of distributed education. ACM Journal of Computer Documentation, 26(4):159–167, November 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Gattis:2002:PIF

[215] Lyn Gattis. Planning and information foraging theories and their value to the novice technical communicator. ACM Journal of Computer Documentation, 26(4):168–175, November 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See commentary [217].

Jack:2002:PIF

[216] Jordynn Jack. Planning and information foraging theories: social implications and extensions. ACM Journal of Computer Documentation, 26(4):

176–180, November 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic).

Eaton:2002:CSI

[217] Angela Eaton. Commentary on “Planning and information foraging theories and their value to the novice technical communicator”. ACM Journal of Computer Documentation, 26(4): 181–183, November 2002. CODEN AJCDBH. ISSN 1527-6805 (print), 1557-9441 (electronic). See [215].

Brockmann:1998:MST

[218] R. John Brockmann. From millwrights to shipwrights to the twenty-first century: explorations in a history of technical communication in the United States. Hampton Press, Cresskill, NJ, USA, 1998. ISBN 1-57273-076-5; 1-57273-077-3. xxii + 464 pp. LCCN T10.5.B75 1998.